# Efficient Error and Variance Estimation

### v4

@@ 4.2 Randomized SVD
The update formula enables efficient algorithms for the leave-one-out error estimator and jackknife variance estimator for the approximation $\boldsymbol{X}$ and derived quantities like projectors onto singular subspaces and truncation of $\boldsymbol{X}$ to rank $r<s$. As one example, Algorithm 2 gives an implementation the randomized SVD with no subspace iteration ($q=0$) with the leave-one-out error estimator $\widehat{\operatorname{Err}}(\boldsymbol{X},\boldsymbol{A})$. The leave-one-out error estimator requires just two lines and runs in $\mathcal{O}(s^{3})$ operations. Further variants are discussed in Section SM1.3 and a MATLAB implementation is provided in LABEL:list:randsvd.
@@ Acknowledgements
We thank Robert Webber for his advice and feedback.

### v5

@@ 4.2 Randomized SVD
The update formula enables efficient algorithms for the leave-one-out error estimator and jackknife variance estimator for the approximation $\boldsymbol{X}$ and derived quantities like projectors onto singular subspaces and truncation of $\boldsymbol{X}$ to rank $r<s$. As one example, Algorithm 2 gives an implementation the randomized SVD with no subspace iteration ($q=0$) with the leave-one-out error estimator $\widehat{\operatorname{Err}}(\boldsymbol{X},\boldsymbol{A})$. The leave-one-out error estimator requires just two lines and runs in $\mathcal{O}(s^{3})$ operations; see Appendix SM5 for a derivation of these two lines. Further variants are discussed in Section SM1.3 and a MATLAB implementation is provided in LABEL:list:randsvd.
@@ Acknowledgements
We thank Per-Gunnar Martinsson, Eitan Levin, and Robert Webber for their advice and feedback.
@@
%% Normalize columns of a possibly complex matrix
function M = cnormc(M)
M = M ./ vecnorm(M,2,1);
end
@@
Section SM5 Derivation of leave-one-out error estimator for randomized SVD
In this section, we provide a derivation for the leave-one-out error estimator for the randomized SVD without subspace iteration $q=0$ given in Algorithm 2 and for $q>0$ given in LABEL:list:randsvd.
SM5.1 The case $q=0$
First suppose $q=0$, and denote $\boldsymbol{G}=(\boldsymbol{R}^{*})^{-1}$ with $j$th column $\boldsymbol{g}_{j}$. By the update formula Eq. 16, we have
$\boldsymbol{Q}^{(j)}\left(\boldsymbol{Q}^{(j)}\right)^{*}=\boldsymbol{Q}\left(\mathbf{I}-\frac{\boldsymbol{g}_{j}^{\vphantom{*}}\boldsymbol{g}_{j}^{*}}{\left\|\boldsymbol{g}_{j}\right\|^{2}}\right)\boldsymbol{Q}^{*}.$
The leave-one-out error estimator is
$\widehat{\operatorname{Err}}^{2}$ $=\frac{1}{s}\sum_{j=1}^{s}\left\|\left(\mathbf{I}-\boldsymbol{Q}^{(j)}\left(\boldsymbol{Q}^{(j)}\right)^{*}\right)\boldsymbol{A}\boldsymbol{\omega}_{j}\right\|^{2}.$
Combining the two previous displays gives
(25) $\widehat{\operatorname{Err}}^{2}=\frac{1}{s}\sum_{j=1}^{s}\left\|\left(\mathbf{I}-\boldsymbol{Q}\boldsymbol{Q}^{*}\right)\boldsymbol{A}\boldsymbol{\omega}_{j}+\frac{\boldsymbol{Q}\boldsymbol{g}_{j}(\boldsymbol{Q}\boldsymbol{g}_{j})^{*}\boldsymbol{A}\boldsymbol{\omega}_{j}}{\left\|\boldsymbol{g}_{j}\right\|^{2}}\right\|^{2}.$
The matrix $\boldsymbol{Q}\boldsymbol{Q}^{*}$ is the orthoprojector onto the column span of $\boldsymbol{A}\boldsymbol{\Omega}$, so $(\mathbf{I}-\boldsymbol{Q}\boldsymbol{Q}^{*})\boldsymbol{A}\boldsymbol{\omega}_{j}=\boldsymbol{0}$. In addition, the Euclidean norm is unitarily invariant, so the leave-one-out estimator simplies as follows:
$\widehat{\operatorname{Err}}^{2}=\frac{1}{s}\sum_{j=1}^{s}\left\|\boldsymbol{g}_{j}\right\|^{-2}\cdot|\boldsymbol{g}_{j}^{*}\boldsymbol{Q}^{*}\boldsymbol{A}\boldsymbol{\omega}_{j}^{\vphantom{*}}|^{2}.$
Since $\boldsymbol{A}\boldsymbol{\Omega}=\boldsymbol{Q}\boldsymbol{R}$, $\boldsymbol{Q}^{*}\boldsymbol{A}\boldsymbol{\omega}_{j}$ is the $j$th column of $\boldsymbol{R}$. By the identity $\boldsymbol{G}=(\boldsymbol{R}^{*})^{-1}=(\boldsymbol{R}^{-1})^{*}$, $\boldsymbol{g}_{j}^{*}$ is the $j$th row of $\boldsymbol{R}^{-1}$. Ergo, $\boldsymbol{g}_{j}^{*}\boldsymbol{Q}^{*}\boldsymbol{A}\boldsymbol{\omega}_{j}^{\vphantom{*}}=1$ and
$\widehat{\operatorname{Err}}^{2}=\frac{1}{s}\sum_{j=1}^{s}\left\|\boldsymbol{g}_{j}\right\|^{-2}.$
This is the formula for the leave-one-out error estimator used in Algorithm 2.
SM5.2 The case $q>0$
Now suppose $q>0$. Following the notation used in LABEL:list:randsvd, we let $\boldsymbol{Z}=\boldsymbol{A}\boldsymbol{\Omega}$ and $\boldsymbol{Y}=(\boldsymbol{A}\boldsymbol{A}^{*})^{q}\boldsymbol{Z}=\boldsymbol{Q}\boldsymbol{R}$. Additionally, let $\boldsymbol{T}$ denote $(\boldsymbol{R}^{*})^{-1}$ with its columns scaled to unit norm. Denote the $j$th column of $\boldsymbol{T}$ and $\boldsymbol{Z}$ as $\boldsymbol{t}_{j}$ and $\boldsymbol{z}_{j}$. Translating Eq. 25 to the present notation, we obtain
$\widehat{\operatorname{Err}}^{2}$ $=\frac{1}{s}\sum_{j=1}^{s}\left\|\left(\mathbf{I}-\boldsymbol{Q}\boldsymbol{Q}^{*}\right)\boldsymbol{z}_{j}+\boldsymbol{Q}\boldsymbol{t}_{j}\boldsymbol{t}_{j}^{*}(\boldsymbol{Q}^{*}\boldsymbol{z}_{j})\right\|^{2}$
$=\frac{1}{s}\left\|\boldsymbol{Z}-\boldsymbol{Q}(\boldsymbol{Q}^{*}\boldsymbol{Z})+\boldsymbol{Q}\boldsymbol{T}\cdot\operatorname{diag}\left(\boldsymbol{t}_{j}^{*}(\boldsymbol{Q}\boldsymbol{z}_{j}):j=1,\ldots,s\right)\right\|_{\rm F}^{2}$
In the second line, we package the sum of squared Euclidean norms of vectors into a squared matrix Frobenius norm. We have obtained the formula that appears in LABEL:list:randsvd.